OpenStudy (anonymous):

an airline allows each passenger 30 kilograms of luggage. Carl has to lower the weight of his suit case by 25% to stay within the limit.

OpenStudy (shinalcantara):

Let 'x' be the original weight of Carl's luggage Since he needs to lower it by 25% to get to the limit which is 30 kg then you'll have it as: x - 0.25x = 30kg ----------------- Using the equation, find x

OpenStudy (anonymous):

how much did the suitcase originally weigh?

OpenStudy (anonymous):

Let M be the maximum weight allowed, and W be the original weight. We know that Carl had to remove 25% of W to make his suitcase's weight be M, and we know that M is 36 kg. (W-25% of W) = M = 36 kg But 25% is just 25/100, so we can write \[W - \left(\begin{matrix}25 \\ 100\end{matrix}\right) W = 36 kg\] multiply both sides by 100 to get rid of the fraction \[100W -25W = 3600kg\] \[75W = 3600kg\] Divide both sides by 75 to find W

OpenStudy (anonymous):

100 * ? = 4800 25 * ? = 1200 4800 - 1200 = 3600 So what is the ?

OpenStudy (anonymous):

48?

OpenStudy (anonymous):

i dont get it -_-

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

i type in 48 and it said it was wrong

OpenStudy (shinalcantara):

@I_Need_Help_With_Home it's actually 30kg in the given

OpenStudy (shinalcantara):

get the value of x in the above equation @jayybee7

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

i tried.. I give up..

OpenStudy (shinalcantara):

x - 0.25x = 30kg

OpenStudy (anonymous):

I dont no how

OpenStudy (shinalcantara):

x - 0.25x = 30kg Simplify 0.75x = 30kg Divide both sides of the equation by 0.75 x = 30/0.75 x = ?

OpenStudy (anonymous):

I still dont get it.. I find it very difficult and frustarting..

OpenStudy (anonymous):

40?

OpenStudy (shinalcantara):

If we let 'x' be the original weight of Carl's luggage and he has to deduct 25% of it to reach the limit which is 30kg, then we can say that what's left is just 75% of the original weight 75% x = 30 changing 75% to decimal you'll have it as 0.75 0.75x = 30 x = 30/0.75 x = 40kg